# A Tutorial on Learning-Based Radio Map Construction: Data, Paradigms, and Physics-Awareness

### 18 Mar 2026 v1

@@ I Introduction
To bypass the constraints of physical data collection, ray tracing (RT) simulators have been widely employed [120, 100]. These computational electromagnetics methods serve as physical engines for synthesizing propagation data. However, the shooting and bouncing rays (SBR) algorithm faces severe scalability bottlenecks [120]. Its computational complexity grows exponentially with the number of interaction orders. This renders ray-optical models too expensive for city-scale dynamic dataset generation [117, 89]. Recent differentiable RT frameworks such as Sionna [48] enable end-to-end gradient computation. They allow joint optimization of simulation parameters and neural network weights. Despite this progress, the millisecond-level latency required by real-time digital twins remains beyond the reach of pure simulation approaches [134].
@@
To address severe degradation under extreme spatial sparsity, the community has transitioned from discriminative point estimation to generative paradigms. Generative adversarial networks (GANs) reformulate RM construction as a conditional generation problem [148, 158, 17, 92, 106, 85]. Diffusion models further decompose generation into a sequence of tractable denoising operations [129, 132, 131, 84, 83, 80, 38, 152, 26]. They offer stable training dynamics and high sample fidelity [47]. Their stochastic nature also enables uncertainty quantification through Monte Carlo sampling over multiple realizations [129, 128].
@@ I-A Related Works
Although RM construction and wireless deep learning have attracted significant attention, existing surveys focus on different aspects. Doha and Abdelhadi [31] provide a comprehensive roadmap for integrating deep learning into wireless receivers. Their scope is confined to temporal and symbol-level processing at the receiver end. This focus is fundamentally distinct from the spatial, environment-scale field reconstruction addressed in this tutorial. Vasudevan and Yuksel [117] offer a broad overview of machine learning for radio propagation. Their coverage emphasizes traditional regression-based algorithms such as random forests and support vector machines. Zeng et al. [146] present a visionary tutorial on the channel knowledge map concept and its network-level applications. Regarding map construction, their discussion relies on classical spatial interpolation techniques like Kriging and tensor completion. Our tutorial, in contrast, treats RM construction as a complex inverse problem and explores advanced generative paradigms. Feng et al. [37] provide a solid categorization of RM estimation methods into model-driven, data-driven, and hybrid approaches. Their architectural coverage extends to ViTs and conditional GANs. However, the existing literature largely views RM construction as a deterministic interpolation or image-to-image translation task. Our tutorial advances beyond this view in three respects. First, we introduce probabilistic generative paradigms to handle extreme spatial sparsity and quantify uncertainty [129, 47]. Second, we systematically review and mathematically adapt optics-inspired neural rendering techniques to the complex-valued RF domain [56, 147, 53]. Third, we propose a rigorous three-level taxonomy for physics-informed integration and formally define the concept of physical hallucinations [132]. A detailed comparison with existing surveys is presented in Table I.
@@
Feature or Dimension Doha et al. (2025) [31] Zeng et al. (2024) [146] Vasudevan et al. (2024) [117] Feng et al. (2025) [35] Our Tutorial (This Work)
@@ I-B Contributions
    We establish a unified architectural taxonomy organized by the forward-inverse problem dichotomy. This taxonomy spans CNNs [69, 66], ViTs [74, 34, 75], GNNs [72, 14, 94], GANs [148, 158, 17], diffusion models [129, 132, 135, 131], and optics-inspired neural rendering methods [56, 55, 147, 53, 26]. Cross-architecture comparisons clarify the trade-offs among inference latency, data efficiency, and spatial modeling capability.
@@ II Preliminaries
Figure 1: The illustration of RM data collection. A comprehensive workflow bridging physical measurements and ray tracing simulations to generate standardized multi-dimensional benchmark datasets.
@@ II-A3 Distinction from Related Concepts
Attribute Radio Map Channel Knowledge Map Radio Environment Map
@@ II-B2 Continuous-Time SDE Generalization
where $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$ is the score function. A score network $\mathbf{s}_{\theta}(\mathbf{x},t)$ approximates this gradient via denoising score matching [119]. The full derivation is provided in Appendix A.
@@ II-B3 Relevance to Wireless Channel Modeling
The continuous-time formulation offers two advantages for wireless applications. First, adaptive ODE/SDE solvers with variable step sizes can accelerate inference [112]. Second, it connects diffusion to stochastic optimal control, enabling physics-informed integration [132]. In practice, most wireless RM methods adopt the discrete DDPM for its simplicity and stable convergence [129, 131, 84, 152].
@@ II-C2 Positional Encoding
where $L$ is the number of frequency bands. In the RF domain, the highest frequency $2^{L-1}\pi$ must resolve variations on the order of $\lambda/2$ [56]. For sub-6 GHz systems where $\lambda\approx 5$1010 cm, the required $L$ is smaller than in optical NeRF. The network is optimized using $L_{2}$ loss: $\mathcal{L}=\sum_{\mathbf{r}\in\mathcal{R}}\|\hat{C}(\mathbf{r})-C_{\mathrm{gt}}(\mathbf{r})\|_{2}^{2}$.
II-C3 Limitations for RF Adaptation
The optical NeRF formulation operates on real-valued RGB intensities under ray-optics assumptions. Wireless RF signals are inherently complex-valued, involving both amplitude and phase [56]. RF propagation at centimeter and millimeter wavelengths is dominated by multipath interference and diffraction [55, 8]. These physical disparities require substantial mathematical redesign, which is detailed in Section VII-A.
II-D 3D Gaussian Splatting
II-D1 Covariance Parameterization
II-D2 Directional Appearance
View-dependent appearance is parameterized with spherical harmonics (SH): $\mathbf{c}(\mathbf{d})=\sum_{l=0}^{L_{\max}}\sum_{m=-l}^{l}\mathbf{k}_{l}^{m}Y_{l}^{m}(\mathbf{d})$. SH functions capture smooth optical illumination well. However, they poorly represent RF angular energy distributions [147, 53]. Coherent multipath at centimeter wavelengths produces sharp angular fluctuations. RF signals are also complex-valued, whereas standard SH encodes real-valued intensities. These issues motivate complex-valued Fourier-Legendre expansions for wireless 3DGS adaptation, as discussed in Section VII-B.
II-D3 Rendering
III Problem Formulation and Taxonomy
Figure 2: Two primary paradigms of neural RM construction. (A) Source-aware modeling, where neural networks act as deterministic surrogates for ray tracing using known environment and transmitter inputs. (B) Source-agnostic reconstruction, representing an ill-posed inverse problem that leverages probabilistic generation to interpolate complete radio fields from sparse, incomplete measurements.
III-A Mathematical Abstractions and Modeling
$C_{E}=C_{geo}+C_{mat}+C_{sem}.$ (13)
@@ III-B2 Transmitter Position Encoding
$\mathbf{P}(j,k)=\exp\!\left(-\frac{(j-a)^{2}+(k-b)^{2}}{2\sigma^{2}}\right),$ (14)
where $(a,b)$ are continuous transmitter coordinates and $\sigma$ controls the spatial spread. This soft representation allows differentiable backpropagation through the location. It is essential for joint RM generation and transmitter localization [133, 154]. The Gaussian heatmap also encodes spatial uncertainty, and its peak sharpens as the model converges.
@@ III-C The Forward Problem: Source-Aware Radio Mapping
Source-aware radio mapping is a forward prediction problem. The objective is to predict global electromagnetic coverage given the propagation environment $\mathbf{E}$ and transmitter parameters $\mathbf{S}$ [69, 129].
Deep neural networks in this setting approximate rigorous computational electromagnetic solvers [37, 18]. They serve as surrogate models for ray tracing engines or full-wave Maxwell’s equation solvers. However, constructing an accurate forward mapping is challenging. The mapping from $\mathbf{E}$ to the signal distribution is a high-sensitivity non-linear function [19]. Minor changes in wall thickness or obstacle displacement can cause large variations due to phase shifts and multipath interference.
@@
The choice of regime depends on the physical coherence conditions [127, 90]. Incoherent addition applies when transmitters operate independently, as in conventional orthogonal frequency division multiple access (OFDMA) inter-cell interference [76, 151]. Each base station transmits asynchronously, and cross-cell signals combine on a power basis. Coherent superposition applies when transmitters share a common phase reference [144, 16]. Examples include distributed multiple-input multiple-output (MIMO) with joint transmission and reconfigurable intelligent surface (RIS)-assisted links. Ignoring phase coherence in such scenarios underestimates the signal-to-noise ratio in constructive regions and overestimates it in destructive zones. This distinction requires architectures that handle global context and complex signal interactions across the spatial domain [74, 34].
@@ III-D The Inverse Problem: Source-Agnostic Radio Mapping
Reconstructing $\mathbf{M}$ from $\mathbf{y}$ is ill-posed because the sampling operator is singular and the solution is non-unique [115, 42]. Multiple source-environment combinations could yield identical sparse observations. The network must incorporate strong structural priors to regularize this inversion [18, 37]. Two canonical priors are widely used. The first is the sparsity of dominant multipath components in the angular domain [15]. Only a limited number of scattering clusters contribute significantly at any location. The second is the low-rank structure of the spatial channel matrix [121]. For narrowband massive MIMO channels, low-rank structure arises from the limited angular spread relative to the array aperture. For spatially distributed RMs, low-rank structure can emerge when the map is organized as a spatial-frequency tensor across multiple sub-bands [141]. All frequency slices share common propagation paths with frequency-invariant delays and angles. However, a single-frequency power map from one transmitter does not inherently exhibit low-rank structure unless additional organization is imposed. Practitioners should verify the applicability of low-rank assumptions for their specific RM modality. The learning model must exploit these priors, embedded within $\mathbf{E}$, to infer the latent source $\mathbf{S}$ [133, 83]. It must also reconstruct physically consistent signal variations in deep shadowing zones where no measurements exist [34, 78].
@@ III-G Discussion on Sampling Paradox
A paradox arises in the source-agnostic inverse problem regarding sampling efficiency. Consider isotropic radiation under the Friis equation: $P_{r}(\mathbf{r}_{i})=P_{t}G_{t}G_{r}\bigl(\frac{\lambda}{4\pi\|\mathbf{r}_{i}-\mathbf{r}_{tx}\|}\bigr)^{2}$. Under free-space propagation, three non-collinear LoS measurements suffice to recover the transmitter location $\mathbf{r}_{tx}$ and power $P_{t}$ via trilateration [8]. However, this three-point sufficiency is only a thought experiment. In practice, multipath fading corrupts LoS measurements with spatial fluctuations of 5–15 dB [40]. This causes substantial errors in trilateration. In urban environments, diffraction, non-line-of-sight (NLoS) blockages, and frequency-selective fading invalidate the geometric inversion entirely [15, 129]. The key insight is that well-placed LoS samples carry redundant geometric information. The true reconstruction challenge lies in NLoS zones where simple distance-power relationships fail. Despite this, many existing works use uniform or random sampling [12, 115]. These strategies oversample predictable LoS regions while undersampling complex NLoS zones. Future strategies should adopt physics-aware adaptive sampling [36, 39]. Sensor placement can be guided by heuristics such as anchoring at diffraction corners, shadow boundaries, and high-entropy zones [19]. Alternatively, active learning frameworks can use predictive uncertainty maps to select high-variance NLoS locations [42]. Such strategies would improve sample efficiency by concentrating measurements where the physical field is least predictable from geometric priors alone.
@@ IV Data Acquisition and Open Datasets
High-quality datasets are essential for data-driven RM construction [146, 117]. This section reviews three data acquisition paradigms: physical measurement, ray tracing simulation, and differentiable simulation. It then discusses preprocessing strategies and catalogs publicly available benchmarks. The section concludes with dataset limitations and selection guidance.
@@ IV-B2 Dominant Path Model
$PL_{total}=20\log_{10}\!\left(\frac{4\pi d}{\lambda}\right)+\sum_{i=1}^{N_{ref}}L_{ref,i}+\sum_{j=1}^{N_{diff}}L_{diff,j}+\sum_{k=1}^{N_{pen}}L_{pen,k}.$ (25)
@@ IV-B3 Differentiable Ray Tracing
IV-C Data Preprocessing and Representation
IV-D Dataset Analysis and Benchmarks
@@
1 RadioMapSeer [69] 2D Path loss 3.5 GHz Sim. (DPM/IRT) 700 maps Dual-engine; 10 m res.; foundational benchmark
@@
3 BART-Lab [72] 2D PL (multi-band) 1.7–5.7 GHz Sim. (Feko/WP) 2,100 maps 5 frequency bands; coarse and fine subsets

### 28 Mar 2026 v4

@@ I Introduction
To bypass the constraints of physical data collection, computational electromagnetic methods including ray tracing (RT) and the dominant path model (DPM) have been widely employed [123, 103]. These computational electromagnetics methods serve as physical engines for synthesizing propagation data. However, the shooting and bouncing rays (SBR) algorithm faces severe scalability bottlenecks [123, 33]. Its computational complexity grows exponentially with the number of interaction orders. This renders ray-optical models too expensive for city-scale dynamic dataset generation [120, 92]. Recent differentiable RT frameworks such as Sionna [50] enable end-to-end gradient computation. They allow joint optimization of simulation parameters and neural network weights. Despite this progress, the millisecond-level latency required by real-time digital twins remains beyond the reach of pure simulation approaches [136].
@@
To address severe degradation under extreme spatial sparsity, the community has transitioned from discriminative point estimation to generative paradigms. Generative adversarial networks (GANs) reformulate RM construction as a conditional generation problem [154, 168, 19, 95, 108, 88]. Diffusion models further decompose generation into a sequence of tractable denoising operations [131, 134, 133, 87, 86, 84, 41, 159, 29]. They offer stable training dynamics and high sample fidelity [49]. Their stochastic nature also enables uncertainty quantification through Monte Carlo sampling over multiple realizations [131].
@@ I-A Related Works
Although RM construction and wireless deep learning have attracted significant attention, existing surveys focus on different aspects. Doha and Abdelhadi [35] provide a comprehensive roadmap for integrating deep learning into wireless receivers. Their scope is confined to temporal and symbol-level processing at the receiver end. This focus is fundamentally distinct from the spatial, environment-scale field reconstruction addressed in this tutorial. Vasudevan and Yuksel [120] offer a broad overview of machine learning for radio propagation. Their coverage emphasizes traditional regression-based algorithms such as random forests and support vector machines. Zeng et al. [152] present a visionary tutorial on the channel knowledge map concept and its network-level applications. Regarding map construction, their discussion relies on classical spatial interpolation techniques like Kriging and tensor completion. Our tutorial, in contrast, treats RM construction as a complex inverse problem and explores advanced generative paradigms. Feng et al. [38] provide a solid categorization of RM estimation methods into model-driven, data-driven, and hybrid approaches. Their architectural coverage extends to ViTs and conditional GANs. However, the existing literature largely views RM construction as a deterministic interpolation or image-to-image translation task. Our tutorial advances beyond this view in three respects. First, we introduce probabilistic generative paradigms to handle extreme spatial sparsity and quantify uncertainty [131, 49]. Second, we systematically review and mathematically adapt optics-inspired neural rendering techniques to the complex-valued RF domain [59, 153]. Third, we propose a rigorous three-level taxonomy for physics-informed integration and formally define the concept of physical hallucinations [134]. A detailed comparison with existing surveys is presented in Table I.
@@
Feature or Dimension Doha and Abdelhadi (2025) [35] Zeng et al. (2024) [152] Vasudevan and Yuksel (2024) [120] Feng et al. (2025) [38] Our Tutorial (This Work)
@@ I-B Contributions
    We establish a unified architectural taxonomy organized by the forward-inverse problem dichotomy. This taxonomy spans CNNs [73, 70], ViTs [78, 37, 79], GNNs [76, 17, 97], GANs [154, 168, 19], diffusion models [131, 134, 137, 133], and optics-inspired neural rendering methods [59, 153, 161, 141, 118, 167]. Cross-architecture comparisons clarify the trade-offs among inference latency, data efficiency, and spatial modeling capability.
@@ II-A1 Electromagnetic Foundation
Attribute Radio Map Channel Knowledge Map Radio Environment Map
@@ II-B2 Continuous-Time SDE Generalization
where $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$ is the score function. A score network $\mathbf{s}_{\theta}(\mathbf{x},t)$ approximates this gradient via denoising score matching [122].
@@ II-B3 Relevance to Wireless Channel Modeling
The continuous-time formulation offers two advantages for wireless applications. First, adaptive ODE/SDE solvers with variable step sizes can accelerate inference [114]. Second, it connects diffusion to stochastic optimal control, enabling physics-informed integration [134]. In practice, many wireless RM methods adopt the discrete DDPM for its simplicity and stable convergence [131, 133, 87, 159].
@@ II-C2 Positional Encoding
where $L$ is the number of frequency bands. In the RF domain, the highest frequency $2^{L-1}\pi$ must resolve variations on the order of $\lambda/2$ [161]. For the upper sub-6 GHz band where $\lambda\approx 5$1010 cm, the required $L$ is smaller than in optical NeRF. The network is optimized using $L_{2}$ loss: $\mathcal{L}=\sum_{\mathbf{r}\in\mathcal{R}}\|\hat{C}(\mathbf{r})-C_{\mathrm{gt}}(\mathbf{r})\|_{2}^{2}$.
@@
II-C3 Limitations for RF Adaptation
The optical NeRF formulation operates on real-valued RGB intensities under ray-optics assumptions. Wireless RF signals are inherently complex-valued, involving both amplitude and phase. RF propagation at centimeter and millimeter wavelengths is dominated by multipath interference and diffraction [58, 11]. These physical disparities require substantial mathematical redesign, which is detailed in Section VII-A.
II-D 3D Gaussian Splatting
II-D1 Covariance Parameterization
II-D2 Directional Appearance
View-dependent appearance is parameterized with spherical harmonics (SH) [64]: $\mathbf{c}(\mathbf{d})=\sum_{l=0}^{L_{\max}}\sum_{m=-l}^{l}\mathbf{k}_{l}^{m}Y_{l}^{m}(\mathbf{d})$. SH functions capture smooth optical illumination well. However, they poorly represent RF angular energy distributions [153, 141]. Coherent multipath at centimeter wavelengths produces sharp angular fluctuations. RF signals are also complex-valued, whereas standard SH encodes real-valued intensities. These issues motivate complex-valued Fourier-Legendre expansions for wireless 3DGS adaptation, as discussed in Section VII-B.
II-D3 Rendering
III Problem Formulation and Taxonomy
III-A Mathematical Abstractions and Modeling
$C_{E}=C_{geo}+C_{mat}+C_{sem}.$ (13)
@@ III-B2 Transmitter Position Encoding
$\mathbf{P}(j,k)=\exp\!\left(-\frac{(j-a)^{2}+(k-b)^{2}}{2\sigma_{s}^{2}}\right),$ (14)
where $(a,b)$ are continuous transmitter coordinates and $\sigma_{s}$ controls the spatial spread. This soft representation allows differentiable backpropagation through the location. It is essential for joint RM generation and transmitter localization [135, 163]. The Gaussian heatmap also encodes spatial uncertainty, and its peak sharpens as the model converges.
@@ III-B3 Condition Embedding Mechanisms
Figure 1: Two primary paradigms of neural RM construction. Part A illustrates source-aware modeling, where neural networks act as deterministic surrogates for ray tracing using known environment and transmitter inputs. Part B illustrates source-agnostic reconstruction, which represents an ill-posed inverse problem that leverages probabilistic generation to interpolate complete radio fields from sparse, incomplete measurements.
@@ III-C The Forward Problem: Source-Aware Radio Mapping
Fig. 1 contrasts the two primary paradigms of neural RM construction: source-aware forward prediction and source-agnostic inverse reconstruction. Source-aware radio mapping is a forward prediction problem. The objective is to predict global electromagnetic coverage given the propagation environment $\mathbf{E}$ and transmitter parameters $\mathbf{S}$ [73, 131].
Deep neural networks in this setting approximate rigorous computational electromagnetic solvers [40, 20]. They serve as surrogate models for ray tracing engines or full-wave Maxwell’s equation solvers. However, constructing an accurate forward mapping is challenging. The mapping from $\mathbf{E}$ to the signal distribution is a highly sensitive nonlinear function [21]. Minor changes in wall thickness or obstacle displacement can cause large variations due to phase shifts and multipath interference.
@@
The choice of regime depends on the physical coherence conditions [130, 93]. Incoherent addition applies when transmitters operate independently, as in conventional orthogonal frequency division multiple access (OFDMA) inter-cell interference [80, 158]. Each base station transmits asynchronously, and cross-cell signals combine on a power basis. Coherent superposition applies when transmitters share a common phase reference [150]. Examples include distributed MIMO with joint transmission and reconfigurable intelligent surface (RIS)-assisted links. Ignoring phase coherence in such scenarios underestimates the signal-to-noise ratio in constructive regions and overestimates it in destructive zones. This distinction requires architectures that handle global context and complex signal interactions across the spatial domain [78, 37].
@@ III-D The Inverse Problem: Source-Agnostic Radio Mapping
Reconstructing $\mathbf{M}$ from $\mathbf{y}$ is ill-posed because the sampling operator is singular and the solution is non-unique [117, 45]. Multiple source-environment combinations could yield identical sparse observations. The network must incorporate strong structural priors to regularize this inversion [20, 40]. Two canonical priors are widely used. The first is the sparsity of dominant multipath components in the angular domain [18]. Only a limited number of scattering clusters contribute significantly at any location. The second is the low-rank structure of the spatial channel matrix [109]. It is important to note that a single-frequency power map from one transmitter does not inherently exhibit low-rank structure. Low-rank structure arises under specific conditions: for narrowband massive MIMO channels, it results from the limited angular spread relative to the array aperture; for spatially distributed RMs, it can emerge when the map is organized as a spatial-frequency tensor across multiple sub-bands where all frequency slices share common propagation paths [144]. Practitioners should verify the applicability of low-rank assumptions for their specific RM modality. The learning model must exploit these priors, embedded within $\mathbf{E}$, to infer the latent source $\mathbf{S}$ [135, 86]. It must also reconstruct physically consistent signal variations in deep shadowing zones where no measurements exist [37, 82].
@@ III-F Target RM and Mapping Function
Figure 2: The overall structure of this tutorial, organized along three intertwined dimensions: data ecosystem, learning paradigms, and physics-awareness.
@@ III-G Discussion on Sampling Paradox
A paradox arises in the source-agnostic inverse problem regarding sampling efficiency. Consider isotropic radiation under the Friis equation: $P_{r}(\mathbf{r}_{i})=P_{t}G_{t}G_{r}\bigl(\frac{\lambda}{4\pi\|\mathbf{r}_{i}-\mathbf{r}_{tx}\|}\bigr)^{2}$. Under free-space propagation in the 2D plane with known antenna height, three non-collinear LoS measurements suffice to recover the transmitter location $\mathbf{r}_{tx}$ and power $P_{t}$ via trilateration [11]. However, this three-point sufficiency is only a thought experiment. In practice, multipath fading corrupts LoS measurements with spatial fluctuations of 5–15 dB [43]. This causes substantial errors in trilateration. In urban environments, diffraction, non-line-of-sight (NLoS) blockages, and frequency-selective fading invalidate the geometric inversion entirely [18, 131]. The key insight is that well-placed LoS samples carry redundant geometric information. The true reconstruction challenge lies in NLoS zones where simple distance-power relationships fail. Despite this, many existing works use uniform or random sampling [15, 117]. These strategies oversample predictable LoS regions while undersampling complex NLoS zones. Future strategies should adopt physics-aware adaptive sampling [39, 42]. Sensor placement can be guided by heuristics such as anchoring at diffraction corners, shadow boundaries, and high-entropy zones [21]. Alternatively, active learning frameworks can use predictive uncertainty maps to select high-variance NLoS locations [45]. Such strategies would improve sample efficiency by concentrating measurements where the physical field is least predictable from geometric priors alone.
@@ IV Data Acquisition and Open Datasets
Figure 3: Illustration of the RM data collection workflow. A comprehensive workflow bridging physical measurements and ray tracing simulations to generate standardized multi-dimensional benchmark datasets.
As outlined in Fig. 2, this tutorial addresses RM construction along three intertwined dimensions: data, paradigms, and physics-awareness. Beginning with the data dimension, high-quality datasets are essential for data-driven RM construction [152, 120]. This section reviews three data acquisition paradigms: physical measurement, ray tracing simulation, and differentiable simulation. It then discusses preprocessing strategies and catalogs publicly available benchmarks. The section concludes with dataset limitations and selection guidance. Fig. 3 illustrates the complete workflow from physical measurements and ray tracing simulations to standardized benchmark datasets.
@@ IV-B2 Dominant Path Model
$PL_{total}$ $=20\log_{10}\!\left(\frac{4\pi d}{\lambda}\right)+\sum_{i=1}^{N_{ref}}L_{ref,i}$
$\hskip 20.00003pt+\sum_{j=1}^{N_{diff}}L_{diff,j}+\sum_{k=1}^{N_{pen}}L_{pen,k}.$ (25)
@@ IV-B3 Differentiable Ray Tracing
1 RadioMapSeer [73] 2D Path loss 5.9 GHz Sim. (DPM/IRT) 56,080 maps Dual-engine; 701 maps $\times$ 80 Tx; 1 m res.; foundational benchmark
@@
3 BART-Lab [76] 2D PL (multi-band) 1.75–5.75 GHz Sim. (Feko/WP) 2,100 maps 5 frequency bands; coarse and fine subsets
@@
8 RadioMap3DSeer [149] 3D Path loss 3.5 GHz Sim. (IRT) 56,080 maps 3D extension of RadioMapSeer
@@
IV-C Data Preprocessing and Representation
IV-D Dataset Analysis and Benchmarks